# Designing Decision Support Systems Using Counterfactual Prediction Sets

### v2

@@ 3 Prediction Sets through a Causal Lens
The experts’ predictions satisfy counterfactual monotonicity if and only if, for any $x\in\mathcal{X}$ and any $\mathcal{C}_{\alpha}$ and $\mathcal{C}_{\alpha^{\prime}}$ such that $Y\in\mathcal{C}_{\alpha}(x)\subseteq\mathcal{C}_{\alpha^{\prime}}(x)$, it holds that $\mathbb{I}\{f_{\hat{Y}}(u,v,\mathcal{C}_{\alpha}(x))=Y\}\geq\mathbb{I}\{f_{\hat{Y}}(u,v,\mathcal{C}_{\alpha^{\prime}}(x))=Y\}$ for any $u\sim P^{\mathcal{M}}(U)$ and $v\sim P^{\mathcal{M}}(V\,|\,X=x)$.
The above assumption directly implies that, for any sample with feature vector $x$ and any $\mathcal{C}_{\alpha}$ and $\mathcal{C}_{\alpha^{\prime}}$ such that $Y\in\mathcal{C}_{\alpha}(x)\subseteq\mathcal{C}_{\alpha^{\prime}}(x)$, if we observe that an expert has succeeded at predicting the ground truth label $Y$ using $\mathcal{C}_{\alpha^{\prime}}$, she would have also succeeded had she used $\mathcal{C}_{\alpha}$ and, conversely, if she has failed at predicting $Y$ using $\mathcal{C}_{\alpha}$, she would have also failed had she used $\mathcal{C}_{\alpha^{\prime}}$, while holding “everything else fixed”. In other words, under the counterfactual monotonicity assumption, the counterfactual dynamics of the expert under certain alternative prediction sets are identifiable and purely deterministic. In what follows, we will leverage this assumption to develop very efficient online algorithms to find the optimal conformal predictor among those using a given calibration set $\mathcal{D}_{\text{cal}}$, i.e., $\alpha^{*}=\mathop{\mathrm{argmax}}_{\alpha}\mathbb{E}_{X,Y,\hat{Y}_{\mathcal{C}_{\alpha}}}[\mathbb{I}\{\hat{Y}_{\mathcal{C}_{\alpha}}=Y\}]$.
Remarks. Since the counterfactual monotonicity assumption lies within level three in the “ladder of causation” [34], we cannot validate it using observational nor interventional experiments. However, the good practical performance of our online algorithms in our human subject study, shown in Figures 2 and 3, suggests that it may hold in the context of the prediction task we have considered. Moreover, in Appendix D, we carry out an additional sensitivity analysis, which shows that the performance of our algorithms degrades gracefully with respect to the amount of violations of the counterfactual monotonicity assumption.
@@ 4 Finding the Optimal Conformal Predictor using Counterfactual Prediction Sets
Given a fixed calibration set $\mathcal{D}_{\text{cal}}=\{(x_{i},y_{i})\}_{i=1}^{m}$, there exist only $m$ different conformal predictors. This is because the empirical quantile $\hat{q}_{\alpha}$, which the subsets $\mathcal{C}_{\alpha}(x_{i})$ depend on, can only take $m$ different values [44]. As a result, to find the optimal conformal predictor, we just need to solve the following maximization problem:
@@
where $\mathcal{A}=\{\alpha_{i}\}_{i\in[m]}$, with $\alpha_{i}=1-i/(m+1)$ and $[m]=\{1,2,\dots,m\}$. However, since we do not know the causal mechanism experts use to make predictions over prediction sets, we need to trade-off exploitation, i.e., maximizing the expected accuracy, and exploration, i.e., learning about the accuracy achieved by the experts under each conformal predictor. To this end, we look at the problem from the perspective of multi-armed bandits [42].
@@
where $\alpha^{*}$ is the optimal parameter value, as defined in Eq. 4. At this point, one could think of resorting to any of the well-known algorithms from the literature on stochastic multi-armed bandits [42], such as UCB1 or successive elimination, to decide which arm to pull, i.e., which $\alpha_{t}$ to use, at each round $t$. These algorithms would achieve an expected regret $\mathbb{E}[R(t)]\leq O(\sqrt{mt\log T})$ for any $t\leq T$, where the expectation is over the randomness in the execution of the algorithms. However, in our problem setting, we can do much better than that—in what follows, we will design an algorithm based on successive elimination that achieves an expected regret $\mathbb{E}[R(t)]\leq O(\sqrt{t\log m\log T})$ for any $t\leq T$.
@@
        $\mathcal{A}_{\text{unexplored}},\gamma,\nu\leftarrow$ Update$(\mathcal{A}_{\text{unexplored}},\gamma,\nu,\tilde{\alpha},x_{t},y_{t},\hat{y}_{t})$Call Algorithm 2}
@@
In the first round, our algorithm pulls the arm whose corresponding parameter value $\tilde{\alpha}$ is the median [^8] The $\frac{m}{2}$-th largest value if $m$ is even or the $\frac{m+1}{2}$-th largest value if $m$ is odd. of all values in $\mathcal{A}_{\text{active}}$. We distinguish two cases. First, assume that the expert has failed at predicting the ground truth label $y_{1}$ using $\mathcal{C}_{\tilde{\alpha}}$, i.e., $\mathbb{I}\{\hat{y}_{1}=y_{1}\}=0$. If $y_{1}\notin C_{\tilde{\alpha}}(x_{1})$, we know that she would have also failed had she used any $\mathcal{C}_{\alpha^{\prime}}$ such that $\alpha^{\prime}>\tilde{\alpha}$ since $\mathcal{C}_{\alpha^{\prime}}(x)\subseteq\mathcal{C}_{\tilde{\alpha}}(x)$. If $y_{1}\in C_{\tilde{\alpha}}(x_{1})$, we know that she would have also failed had she used any $\mathcal{C}_{\alpha^{\prime}}$ such that $\alpha^{\prime}<\tilde{\alpha}$ due to the counterfactual monotonicity assumption. Second, assume that the expert has succeeded at predicting $y_{1}$ using $\mathcal{C}_{\tilde{\alpha}}$, i.e., $\mathbb{I}\{\hat{y}_{1}=y_{1}\}=1$. Then, for any $\alpha^{\prime}>\tilde{\alpha}$, we know that, if $y_{1}\in\mathcal{C}_{\alpha^{\prime}}(x_{1})$, the same expert would have also succeeded had she used $\mathcal{C}_{\alpha^{\prime}}$ due to the counterfactual monotonicity assumption and, if $y_{1}\notin\mathcal{C}_{\alpha^{\prime}}(x_{1})$, the expert would have trivially failed had she used $\mathcal{C}_{\alpha^{\prime}}$. In both cases, the algorithm observes the reward for one arm and counterfactually infers the reward for at least one half [^9] One half is $\frac{m}{2}$ or $\frac{m}{2}-1$ values if $m$ is even and $\frac{m-1}{2}$ values if $m$ is odd. of the arms in $\mathcal{A}_{\text{active}}$. In the next $O(\log m)$ rounds, it repeats the same reasoning, it pulls the arm whose parameter value is the median of the remaining (at most) half of the arms whose reward has not yet observed or counterfactually inferred, until it has observed or counterfactually inferred the reward of all arms at least once. Then, it computes
@@
[…]
@@
Remarks. Counterfactual SE assumes that, every time an expert predicts a sample, the ground truth label $y$ is (eventually) observed. However, there may be scenarios in which the ground truth labels are only observed during the design phase of the decision support system, but not during deployment. In those cases, it may be more meaningful to look at the problem from the perspective of best-arm identification and derive theoretical guarantees regarding the probability that Counterfactual SE chooses the best arm [42], which is left as future work. In this context, it is worth noting that SE and UCB1 offer desirable theoretical guarantees for best-arm identification [17, 2].
@@ 5 Evaluation via Human Subject Study
In this section, we conduct a large-scale human subject study and show that: a) human experts are more likely to succeed at predicting the ground truth label under smaller prediction sets, providing strong evidence that the interventional monotonicity assumption (Assumption 1) holds; b) counterfactual successive elimination and counterfactual UCB1 achieve a significant improvement in expected regret in comparison with their vanilla implementations; c) a strict implementation of our decision support system, which adaptively limits experts’ level of agency, offers greater performance than a lenient implementation, which allows experts to always exercise their own agency.101010 All experiments ran on a Mac OS machine with an M1 processor and $16$GB memory.
Human subject study setup. We gathered $194{,}407$ label predictions from $2{,}751$ human participants for $1{,}200$ unique images from the ImageNet16H dataset [43] using Prolific. Our experimental protocol received approval from the Institutional Review Board (IRB) at the University of Saarland, each participant was rewarded with $9$£ per hour pro-rated, following Prolific’s payment principles, and consented to participate by filling a consent form that included a detailed description of the study processes, and the collected data did not include any personally identifiable information. Each image in the ImageNet16H dataset belongs to one of $16$ different categories (e.g., animals, vehicles as well as every day objects), which serve as labels. In our study, we used always the same classifier, namely the pre-trained VGG-19 [41] after $10$ epochs of fine-tuning as provided by Steyvers et al. [43] and a fixed calibration set of $120$ images, picked at random. [^11] The model and the dataset ImageNet16H are publicly available at https://osf.io/2ntrf/. The average accuracy of the classifier (over the images not in the calibration set) is $0.848$. For each image, we first computed all possible prediction sets that any conformal predictor using the above classifier and calibration set could construct. Then, we created $715$ questionnaires, each with a set of images and, for each image, a multiple choice question using a prediction set. Under the strict implementation of our system, the multiple choice options included only the label values of the corresponding prediction set and, under the lenient implementation, they additionally included an option “Other”, which allowed participants to pick a label value outside the prediction set. Under both implementations, the questionnaires covered all possible pairs of images and prediction sets. We provide additional information about the ImageNet16H dataset, further implementation details as well as screenshots of the questionnaires and the consent form in Appendix B. In what follows, if not said otherwise, the results refer to the strict implementation of our system.
@@
Figure 3: Empirical success probability achieved by all experts across all images using the strict and lenient implementation of our system $\mathcal{C}_{\alpha}$ with different $\alpha$ values. For the strict implementation, we annotate the optimal $\alpha$ value, the $\alpha$ values found by the algorithms by [44] and by counterfactual UCB1, as well as the average success probability achieved by the set of $\alpha$ values that remain active after running counterfactual SE. For the lenient implementation, we annotate the $\alpha$ value used by [3]. The average accuracy of the classifier used by both the strict and the lenient implementation of our system is $0.848$ and the empirical success probability achieved by the experts on their own is $0.760$. The shaded areas correspond to a $95\%$ confidence interval.
@@
Strict vs. lenient implementation of our system. One of the key motivations to forcefully ask experts to predict label values from the prediction sets provided by conformal prediction is to be able to trade-off how frequently the system misleads experts and the difficulty of the task the expert needs to solve [44]. However, one could argue that a more lenient system, which allows experts to predict label values from outside the prediction sets as suggested by [3], may offer greater performance since, in principle, it does not forcefully mislead an expert whenever the prediction set does not contain the ground truth label. Here, we provide empirical evidence that suggests that this is not the case. Figure 3 demonstrates that a strict implementation of our system consistently offers greater performance than a lenient implementation across the full spectrum of competitive $\alpha$ values, [^14] We have excluded values of $\alpha>0.5$ to improve visibility, however, we include the full figure in Appendix E. where we allow experts to predict any label value from $\mathcal{Y}$ whenever a prediction set is empty under both the strict and the lenient implementation of our system. In Appendix E, we further investigate why the strict implementation is superior. Figure 3 also demonstrates that counterfactual UCB1 and counterfactual SE offer a significant advantage over the algorithm proposed by [44].

### v3

@@ 3 Prediction Sets through a Causal Lens
The experts’ predictions satisfy counterfactual monotonicity if and only if, for any $x\in\mathcal{X}$ and any $\mathcal{C}_{\alpha}$ and $\mathcal{C}_{\alpha^{\prime}}$ such that $Y\in\mathcal{C}_{\alpha}(x)\subseteq\mathcal{C}_{\alpha^{\prime}}(x)$, it holds that
$\mathbb{I}\{f_{\hat{Y}}(u,v,\mathcal{C}_{\alpha}(x))=Y\}\geq\mathbb{I}\{f_{\hat{Y}}(u,v,\mathcal{C}_{\alpha^{\prime}}(x))=Y\}$
for any $u\sim P^{\mathcal{M}}(U)$ and $v\sim P^{\mathcal{M}}(V\,|\,X=x)$.
The above assumption is a sufficient condition for interventional monotonicity since we can obtain the interventional monotonicity inequality given by Eq. 3, by taking the expectation in both sides of the counterfactual monotonicity inequality, i.e., $\mathbb{I}\{f_{\hat{Y}}(u,v,\mathcal{C}_{\alpha}(x))=Y\}\geq\mathbb{I}\{f_{\hat{Y}}(u,v,\mathcal{C}_{\alpha^{\prime}}(x))=Y\}$ over $u,v$. Further, the counterfactual monotonicity assumption directly implies that, for any sample with feature vector $x$ and any $\mathcal{C}_{\alpha}$ and $\mathcal{C}_{\alpha^{\prime}}$ such that $Y\in\mathcal{C}_{\alpha}(x)\subseteq\mathcal{C}_{\alpha^{\prime}}(x)$, if we observe that an expert has succeeded at predicting the ground truth label $Y$ using $\mathcal{C}_{\alpha^{\prime}}$, she would have also succeeded had she used $\mathcal{C}_{\alpha}$ and, conversely, if she has failed at predicting $Y$ using $\mathcal{C}_{\alpha}$, she would have also failed had she used $\mathcal{C}_{\alpha^{\prime}}$, while holding “everything else fixed”. In other words, under the counterfactual monotonicity assumption, the counterfactual dynamics of the expert under certain alternative prediction sets are identifiable and purely deterministic. However, note that this does not prevent the factual dynamics of the expert from being stochastic and fallible, as extensively argued in psychology and behavior literature (Kahneman and Tversky,, 1973; Tversky and Kahneman,, 1981; Kahneman and Tversky,, 1984; Loewenstein and Prelec,, 1992; Loewenstein et al.,, 2003).
In what follows, we will leverage this assumption to develop very efficient online algorithms to find the optimal conformal predictor among those using a given calibration set $\mathcal{D}_{\text{cal}}$, i.e., $\alpha^{*}=\mathop{\mathrm{argmax}}_{\alpha}\mathbb{E}_{X,Y,\hat{Y}_{\mathcal{C}_{\alpha}}}[\mathbb{I}\{\hat{Y}_{\mathcal{C}_{\alpha}}=Y\}]$.
Remarks. Since the counterfactual monotonicity assumption lies within level three in the “ladder of causation” (Pearl,, 2009), we cannot validate it using observational nor interventional experiments. However, the good practical performance of our online algorithms in our human subject study, shown in Figures 2 and 3, suggests that it may hold in the context of the prediction task we have considered. Moreover, in Appendix D, we carry out an additional sensitivity analysis, which shows that the performance of our algorithms degrades gracefully with respect to the amount of violations of the counterfactual monotonicity assumption.
@@ 4 Finding the Optimal Conformal Predictor using Counterfactual Prediction Sets
Given a fixed calibration set $\mathcal{D}_{\text{cal}}=\{(x_{i},y_{i})\}_{i=1}^{m}$, there exist only $m$ different conformal predictors. This is because the empirical quantile $\hat{q}_{\alpha}$, which the subsets $\mathcal{C}_{\alpha}(x_{i})$ depend on, can only take $m$ different values (Straitouri et al.,, 2023). As a result, to find the optimal conformal predictor, we just need to solve the following maximization problem:
@@
where $\mathcal{A}=\{\alpha_{i}\}_{i\in[m]}$, with $\alpha_{i}=1-i/(m+1)$ and $[m]=\{1,2,\dots,m\}$. However, since we do not know the causal mechanism experts use to make predictions over prediction sets, we need to trade-off exploitation, i.e., maximizing the expected accuracy, and exploration, i.e., learning about the accuracy achieved by the experts under each conformal predictor. To this end, we look at the problem from the perspective of multi-armed bandits (Slivkins,, 2019).
@@
where $\alpha^{*}$ is the optimal parameter value, as defined in Eq. 4. At this point, one could think of resorting to any of the well-known algorithms from the literature on stochastic multi-armed bandits (Slivkins,, 2019), such as UCB1 or successive elimination, to decide which arm to pull, i.e., which $\alpha_{t}$ to use, at each round $t$. These algorithms would achieve an expected regret $\mathbb{E}[R(t)]\leq O(\sqrt{mt\log T})$ for any $t\leq T$, where the expectation is over the randomness in the execution of the algorithms. However, in our problem setting, we can do much better than that—in what follows, we will design an algorithm based on successive elimination that achieves an expected regret $\mathbb{E}[R(t)]\leq O(\sqrt{t\log m\log T})$ for any $t\leq T$.
@@
        $\mathcal{A}_{\text{unexplored}},\gamma,\nu\leftarrow$ Alg.2$(\mathcal{A}_{\text{unexplored}},\gamma,\nu,\tilde{\alpha},x_{t},y_{t},\hat{y}_{t})$
@@
In the first round, our algorithm pulls the arm whose corresponding parameter value $\tilde{\alpha}$ is the median [^8] The $\frac{m}{2}$-th largest value if $m$ is even or the $\frac{m+1}{2}$-th largest value if $m$ is odd. of all values in $\mathcal{A}_{\text{active}}$. We distinguish two cases. First, assume that the expert has failed at predicting the ground truth label $y_{1}$ using $\mathcal{C}_{\tilde{\alpha}}$, i.e., $\mathbb{I}\{\hat{y}_{1}=y_{1}\}=0$. If $y_{1}\notin C_{\tilde{\alpha}}(x_{1})$, we know that she would have also failed had she used any $\mathcal{C}_{\alpha^{\prime}}$ such that $\alpha^{\prime}>\tilde{\alpha}$ since $\mathcal{C}_{\alpha^{\prime}}(x)\subseteq\mathcal{C}_{\tilde{\alpha}}(x)$. If $y_{1}\in C_{\tilde{\alpha}}(x_{1})$, we know that she would have also failed had she used any $\mathcal{C}_{\alpha^{\prime}}$ such that $\alpha^{\prime}<\tilde{\alpha}$ due to the counterfactual monotonicity assumption. Second, assume that the expert has succeeded at predicting $y_{1}$ using $\mathcal{C}_{\tilde{\alpha}}$, i.e., $\mathbb{I}\{\hat{y}_{1}=y_{1}\}=1$. Then, for any $\alpha^{\prime}>\tilde{\alpha}$, we know that, if $y_{1}\in\mathcal{C}_{\alpha^{\prime}}(x_{1})$, the same expert would have also succeeded had she used $\mathcal{C}_{\alpha^{\prime}}$ due to the counterfactual monotonicity assumption and, if $y_{1}\notin\mathcal{C}_{\alpha^{\prime}}(x_{1})$, the expert would have trivially failed had she used $\mathcal{C}_{\alpha^{\prime}}$. In both cases, the algorithm observes the reward for one arm and counterfactually infers the reward for at least one half [^9] One half is $\frac{m}{2}$ or $\frac{m}{2}-1$ values if $m$ is even and $\frac{m-1}{2}$ values if $m$ is odd. of the arms in $\mathcal{A}_{\text{active}}$.
In the next $O(\log m)$ rounds, it repeats the same reasoning, it pulls the arm whose parameter value is the median of the remaining (at most) half of the arms whose reward has not yet observed or counterfactually inferred, until it has observed or counterfactually inferred the reward of all arms at least once. Then, it computes
@@
Given a calibration set $\mathcal{D}_{\text{cal}}=\{(x_{i},y_{i})\}_{i=1}^{m}$ and a maximum number of rounds $T\geq\sqrt{m}$, Counterfactual SE is guaranteed to achieve expected regret $\mathbb{E}[R(t)]\leq O\left(\sqrt{t\log m\log T}\right)$ for any $t\leq T$.
@@
Remarks. Counterfactual SE assumes that, every time an expert predicts a sample, the ground truth label $y$ is (eventually) observed. However, there may be scenarios in which the ground truth labels are only observed during the design phase of the decision support system, but not during deployment. In those cases, it may be more meaningful to look at the problem from the perspective of best-arm identification and derive theoretical guarantees regarding the probability that Counterfactual SE chooses the best arm (Slivkins,, 2019), which is left as future work. In this context, it is worth noting that SE and UCB1 offer desirable theoretical guarantees for best-arm identification (Even-Dar et al.,, 2006; Audibert et al.,, 2010).
@@ 5 Evaluation via Human Subject Study
In this section, we conduct a large-scale human subject study and show that:
  • a)
    human experts are more likely to succeed at predicting the ground truth label under smaller prediction sets, providing strong evidence that the interventional monotonicity assumption (Assumption 1) holds;
  • b)
    counterfactual successive elimination and counterfactual UCB1 achieve a significant improvement in expected regret in comparison with their vanilla implementations;
  • c)
    a strict implementation of our decision support system, which adaptively limits experts’ level of agency, offers greater performance than a lenient implementation, which allows experts to always exercise their own agency. [^10] All experiments ran on a Mac OS machine with an M1 processor and $16$GB memory.
Human subject study setup. To construct our dataset ImageNet16H-PS, we gathered $194{,}407$ label predictions from $2{,}751$ human participants for $1{,}200$ unique images from the ImageNet16H dataset (Steyvers et al.,, 2022) using Prolific. Our experimental protocol received approval from the Institutional Review Board (IRB) at the University of Saarland, each participant was rewarded with $9$£ per hour pro-rated, following Prolific’s payment principles, and consented to participate by filling a consent form that included a detailed description of the study processes, and the collected data did not include any personally identifiable information. Each image in the ImageNet16H dataset belongs to one of $16$ different categories (e.g., animals, vehicles as well as every day objects), which serve as labels. In our study, we used always the same classifier, namely the pre-trained VGG-19 (Simonyan and Zisserman,, 2015) after $10$ epochs of fine-tuning as provided by Steyvers et al., (2022) and a fixed calibration set of $120$ images, picked at random. [^11] The model and the dataset ImageNet16H are publicly available at https://osf.io/2ntrf/. The average accuracy of the classifier (over the images not in the calibration set) is $0.848$. For each image, we first computed all possible prediction sets that any conformal predictor using the above classifier and calibration set could construct. Then, we created $715$ questionnaires, each with a set of images and, for each image, a multiple choice question using a prediction set. Under the strict implementation of our system, the multiple choice options included only the label values of the corresponding prediction set and, under the lenient implementation, they additionally included an option “Other”, which allowed participants to pick a label value outside the prediction set. Under both implementations, the questionnaires covered all possible pairs of images and prediction sets. We provide additional information about the ImageNet16H dataset, further implementation details as well as screenshots of the questionnaires and the consent form in Appendix B. In what follows, if not said otherwise, the results refer to the strict implementation of our system.
@@
Figure 3: Empirical success probability achieved by all experts across all images using the strict and lenient implementation of our system $\mathcal{C}_{\alpha}$ with different $\alpha$ values. For the strict implementation, we annotate the optimal $\alpha$ value, the $\alpha$ values found by the algorithms by Straitouri et al., (2023) [1] and by counterfactual UCB1, as well as the average success probability achieved by the set of $\alpha$ values that remain active after running counterfactual SE. For the lenient implementation, we annotate the $\alpha$ value used by Babbar et al., (2022) [2]. The average accuracy of the classifier used by both the strict and the lenient implementation of our system is $0.848$ and the empirical success probability achieved by the experts on their own is $0.760$. The shaded areas correspond to a $95\%$ confidence interval.
Strict vs. lenient implementation of our system. One of the key motivations to forcefully ask experts to predict label values from the prediction sets provided by conformal prediction is to be able to trade-off how frequently the system misleads experts and the difficulty of the task the expert needs to solve (Straitouri et al.,, 2023). However, one could argue that a more lenient system, which allows experts to predict label values from outside the prediction sets as suggested by Babbar et al., (2022), may offer greater performance since, in principle, it does not forcefully mislead an expert whenever the prediction set does not contain the ground truth label. Here, we provide empirical evidence that suggests that this is not the case. Figure 3 demonstrates that a strict implementation of our system consistently offers greater performance than a lenient implementation across the full spectrum of competitive $\alpha$ values, [^14] We have excluded values of $\alpha>0.5$ to improve visibility, however, we include the full figure in Appendix E. where we allow experts to predict any label value from $\mathcal{Y}$ whenever a prediction set is empty under both the strict and the lenient implementation of our system. In Appendix E, we further investigate why the strict implementation is superior. Figure 3 also demonstrates that counterfactual UCB1 and counterfactual SE offer a significant advantage over the algorithm proposed by Straitouri et al., (2023), which uses a stylized expert model.